OpenStudy (anonymous):

y varies directly as the square of R. If y = 256 when R = 8, find y when R = 6.

OpenStudy (anonymous):

Y = K R^2 or K = Y/R^2 Becasuse y =256 when R=8 means that K = 256/64 =4 so Y = 4R^2 Y = 4(6)^2 = 4*36 =144 when R=6